OpenStudy (anonymous):

Find the fifth roots of the complex number. cos(12pi/7)+isin(12pi/7) The fifth root of the smallest argument is __? form a+bi

OpenStudy (amistre64):

demorgans

OpenStudy (amistre64):

\[r^5\sqrt{cos(5t)+sin(5t)}\] if i remember it right

OpenStudy (earthcitizen):

From Euler's formula \[re ^{i \theta}=\cos(\theta)+isin(\theta)\] \[r=1, \theta=12\pi/7\] therefore\[e ^{i(12\pi/7)}= \cos(12\pi/7) +isin(12\pi/7)\]

OpenStudy (amistre64):

5th root of the smallest arguement?

OpenStudy (anonymous):

yes, thats what it's asking for.

OpenStudy (amistre64):

demovers is what i was thinking; and 5 would be 1/5

OpenStudy (earthcitizen):

using de Moivre's theorem \[z ^{n}=(re ^{i \theta})^{n}\] \[n=5, \theta=12\pi/7\] \[\therefore z ^{5}=1^{5}e ^{i(5(12\pi/7))}= 1^{5}(\cos(5(12\pi/7))+isin(5(12\pi/7))\]

OpenStudy (anonymous):

Please refer my ans on your earlier post

OpenStudy (earthcitizen):

however, polar form can be used where\[r ^{1/n}<(\theta+360k)/n\]

OpenStudy (amistre64):

12pi/35 * 180/pi 12*180/35 =abt 60 cos(60) is smaller than sin(60) but i got no idea if thats a good answer :/

OpenStudy (earthcitizen):

\[\theta=12\pi/7 = 308.6^{o}, r= 1, n=5, k=0\] \[@ k =0 \therefore z ^{5}= 1^{1/5}(308.6/5)=1<61.72^{o}\] \[@ k =1\therefore z ^{5}= 1<(308.6+360)/5= 1<133.72^{0}\] \[...@k = 4, z=1<(308.6+(360\times4))/5)=1<349.72^{?}\] convert back to rectangular form